# What’s Wrong with Your Code Generated by Large Language Models? An Extensive Study

### v1

@@ 3. LLM-Based Code Generation
3.1. Effectiveness of LLMs in Code Generation.
@@
Within the constraints of our experiment, we consider the generated code to be correct if it passes compiler verification, executes without error, and outputs the expected result for predefined inputs. Additionally, correct code must pass all unit test cases in the benchmark, as different unit test cases cover various semantics. Only by passing all tests can the generated code demonstrate complete fulfillment of all the requirements.
Based on Table 2, we can observe that GPT-4 performs the best, followed by Claude-3, then GPT-3.5, with all open-source models performing worse than GPT-3.5. Specifically in accordance with HumanEval+ and MBPP+, GPT-4 demonstrates the best performance on HumanEval+, while Claude-3 performs best on MBPP+. For APPS+, GPT-4 maintains a 71.5% passing rate on simpler Introductory problems. However, when faced with more complex problems (i.e., Interview and Competition), the accuracy of all models drops significantly. GPT-3.5 and all open-source models have less than a 10% passing rate on Competition problems.
Finding 1: Closed-source LLMs outperform open-source LLMs, particularly when handling more complex tasks. Specifically, GPT-4 performs with the best accuracy across different benchmarks.
Additionally, we investigated the token count of problem descriptions in the three benchmarks, as well as the lines of code, cyclomatic complexity, and number of APIs used in the corresponding canonical solutions. The token count of problem descriptions and the aforementioned code characteristics can, to some extent, represent the complexity of the problems. The HumanEval+ and MBPP+ have concise problem descriptions, short code lengths, low complexity, and fewer API calls. In contrast, APPS+ is more complex, with extensive problem descriptions and intricate code, including 447.6 tokens in the problem description, 23.8 lines of code, 5.5 cyclomatic complexity and 13 API calls. Based on the manual review of APPS+, we found that the problem descriptions contain an amount of redundant and irrelevant information, such as background knowledge unrelated to algorithms. Therefore, as the length and complexity of the problems increase, the accuracy of the code generated by the models significantly decreases, indicating that the difficulty and characteristics of various tasks have a substantial impact on model capabilities.
Finding 2: As the complexity of the benchmarks increases, the accuracy of LLMs tends to decline, especially when the lines of code and the number of APIs increase.
@@
Figure 1. The code characteristic differences between code generated by models and canonical solutions on HumanEval+. $C$ denotes cyclomatic complexity.
Figure 2. The difference of comment numbers between correct and incorrect code generated by LLMs. SC2 denotes StarCoder-2, DC denotes DeepSeekCoder, LL3 denotes Llama-3, and CL3 denotes Claude-3.
Figure 3. Taxonomy of bugs that occurred in code generated by LLMs.
3.2. Factors for LLMs in Code Generation
@@
In this section, we analyze specific characteristics in the generated code, including the number of lines of code, cyclomatic complexity, number of APIs, and the number of comment lines.
@@
First, we investigated the differences in code characteristics between the correct generated code and the canonical solution. Figure 1 shows the differences in the number of lines of code, cyclomatic complexity, and number of APIs on HumanEval+. It can be observed that the code generated by the models typically has fewer lines but slightly higher complexity, while the number of APIs is similar to that of the canonical solutions. Notably, code generated by Claude-3 tends to include fewer lines of code and lower complexity while utilizing more APIs, which is likely due to Claude-3 utilizing APIs to replace parts of the complex code implementation. Due to space limitations in this paper, we place the experimental results for the remaining benchmarks in the open-source repository (ar2, 2023).
@@
Finding 3: The code generated by the models typically has fewer lines but higher code complexity and a similar API number as compared to the canonical solutions.
Furthermore, we observed a distribution pattern in the number of comments in correct code versus incorrect code, as shown in Figure 2. Currently, researchers have found that introducing a thinking process during the inference or supervised fine-tuning phase can increase the reasoning and understanding ability of LLMs, such as in math reasoning (Imani et al., 2023; Wang et al., 2022a; Vacareanu et al., 2024) and task planning (Wei et al., 2022; Sharan et al., 2023). Adding comments in the code corpus can also enhance the performance of LLMs in code-relevant tasks (Yin et al., 2018; Ahmad et al., 2021; Wang et al., 2021; Neelakantan et al., 2022). In code generation, code comments generated by LLMs can enhance their understanding of code problems and help them better generate the subsequent code (Song et al., 2024; Muennighoff et al., [n. d.]). However, our experiment revealed that the variability in the number of comments differs between correct and incorrect code, with incorrect code often showing a wider spread. Figure 2 illustrates the number of comments in correct and incorrect code for different LLMs across various benchmarks. For HumanEval+, incorrect code, particularly from Claude-3 and Deepseek-Coder, tends to have more and more varied comments than correct code. Similarly, in the APPS+ dataset, additional models tend to include comments in both correct and incorrect code, with a tendency to produce a higher median number of comments in incorrect code. This phenomenon could be attributed to two potential factors. Firstly, the complex problems present in the code corpora used to train these LLMs often contain more comments. Consequently, LLMs may learn to replicate this pattern, resulting in the generation of more comments when tackling complex challenges. Secondly, LLMs might generate additional comments as a form of contextual or explanatory content when they are uncertain about the correctness of the generated code. However, these comments do not significantly improve the LLMs’ accuracy, as the correctness rate remains low. Thus, the incorrect code contains more comments than the correct code.
Finding 4: Incorrect code is generally associated with a higher number of comments, suggesting that LLMs might include more explanatory comments when the code is complex and less likely to be correct.
4. Bugs in Code Generated by LLMs
@@
To fully understand the root causes of bugs in experiments, we categorized 5,072 bugs generated from seven LLMs using script-based initial annotation and manual detailed annotation, identifying three primary bug types and 12 secondary bug types (Section 4.1). We further performed a comprehensive bug analysis, including an investigation of the distribution of various bugs, and discussed findings for different bug types (Section 4.2).
4.1. Taxonomy of Code Generation Bugs
4.1.1. Methodology.
@@ 4.1. Taxonomy of Code Generation Bugs
Inspired by existing research, we first annotated all bugs into three primary bug types: Syntax Bug, Runtime Bug, and Functional Bug. These three primary bug types, identified using the Python interpreter’s output, encompass all possible bug types. Based on the primary bug types above, we adopted a two-stage process involving script-based initial annotation and manual detailed annotation. The initial script-based annotation aims to efficiently annotate the primary bug types using output from the Python interpreter, enhancing the manual detailed annotation. The manual detailed annotation involved experts discussing and refining the taxonomy of secondary bug types, ensuring that all bugs were correctly annotated into appropriate secondary types.
@@
In the first stage, the script collected output from the Python interpreter during code execution and used regular expressions for primary bug annotation. Syntax bugs are identified by matching “Syntax Invalid” or “Syntax Error” in the error messages. Runtime bugs are detected by matching the “Traceback” field. Functional bugs are recognized by the “AssertionError” phrase, as assert statements are used to verify outputs in the experiment. It should be noted that the annotation priority is: syntax bug, runtime bug, and functional bug. This means we only identify a functional bug if the code does not exhibit syntax or runtime bugs. However, the interpreter does not provide detailed error information for “AssertionError”. Therefore, the script utilized GPT-4 to generate three possible root causes along with explanations, aiding in subsequent manual analysis. We also discuss the accuracy of three annotation methods (i.e., fully automated, fully manual, and AI-assisted) for labeling bug types in Section 7.
@@
In the second stage, 22 experts with a background in software engineering, reviewed the incorrect code with the primary bug types and independently created a taxonomy for secondary bug types. Then, the experts combined each individual taxonomy and discussed their findings to establish a final taxonomy. Finally, different experts double-checked the annotation to ensure all bugs were correctly annotated. For bugs with disagreements in annotation, all experts convened for a joint discussion until a consensus was reached. Additionally, experts were required to review and revise the annotation of similar bugs to ensure consistency. Notably, because the code generation process is based on “next code token prediction”, the generated code directly influences subsequent code generation. Therefore, we annotated bugs based on the root cause of the first erroneous code in the generated code. Additionally, during the code annotation process, we found that some problems have incomplete or ambiguous functionality descriptions. The models attempt to infer the missing details to complete the code generation, leading to failures in unit tests. We refer to these issues as Ambiguous Problems. The two stages collectively took 2,400 person-hours and culminated in the classification of a taxonomy of three primary bug types and 12 secondary bug types, as shown in Figure 3. Detailed introductions of each bug type will be discussed in the following sections.
4.1.2. Type A: Syntax Bug.
@@
Syntax bugs violate the grammatical rules of the programming language being used. These bugs are detected by the Python interpreter when it tries to parse before the code execution. There are three secondary syntax bug types: Incomplete Syntax Structure, Incorrect Indentation, and Library Import Error.
@@
A.1 Incomplete Syntax Structure. An incomplete syntax structure indicates that the generated code includes an open or partially written syntax element that has not been properly completed. This type of bug includes incomplete statements, unmatched parentheses, un-closed quotes, or missing colons.
print(" ".join(map(str, range(1, n + 1))) # unclosed Parentheses
@@
A.2 Incorrect Indentation Python relies on indentation to define the scope of loops, conditionals, functions, and other code blocks. Incorrect indentation violates the syntax structure, causing syntax errors or inconsistent semantics. As shown in the following example, the indentation level of the second line does not match the subsequent lines within the same code block, which violates the Python specification.
sides = [a, b, c] # inconsistent indentation level
sides.sort()
@@
A.3 Library Import Error. Library import allows Python programs to utilize external code, avoiding redundant development. Common import errors include missing import statements and incorrect import levels. In the following example, the generated code incorrectly imports all public functions from the heapq library within the function body. However, this operation is only permitted outside the function body.
from heapq import * # import * only allowed at module level
4.1.3. Type B: Runtime Bug.
@@
Runtime Bugs refer to bugs that arise when the code fails to conform to the runtime specification, which is detected during execution. Based on the taxonomy, there are five secondary runtime bugs: API Misuse, Definition Missing, Incorrect Boundary Condition Check, Incorrect Argument, and Minors.
@@
tup.sort() # ’tuple’ object has no attribute ’sort
Table 3. Types of bugs introduced during code translation by seven popular LLMs on three widely used benchmarks (i.e., HumanEval+, MBPP+, and APPS+). SC2 denotes StarCoder-2, DC denotes DeepSeekCoder, LL3 denotes Llama-3, and CL3 denotes Claude-3. All values are in %.
@@
Open-Source Closed-Source Open-Source Closed-Source Open-Source Closed-Source
SC2 DC LL3 Phi3 GPT4 GPT3.5 CL3 SC2 DC LL3 Phi3 GPT4 GPT3.5 CL3 SC2 DC LL3 Phi3 GPT4 GPT3.5 CL3
A.1 Incomplete Syntax Structure 0.0 0.0 0.0 0.0 0.0 0.6 0.0 0.3 0.5 2.4 1.9 0.0 0.5 0.0 0.3 3.5 2.7 0.8 0.2 1.5 0.2
A.2 Incorrect Indentation 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.2 0.0 0.0 0.0 1.2 0.0
A.3 Library Import Error 3.7 2.4 10.4 0.0 0.6 2.4 0.0 0.3 0.0 0.8 0.0 0.3 0.5 0.0 0.0 0.0 0.5 0.2 1.0 2.0 2.5
A Syntax Bug 3.7 2.4 10.4 0.0 0.6 3.0 0.0 0.5 0.5 3.2 1.9 0.3 1.0 0.0 0.3 3.7 3.2 1.0 1.2 4.7 2.7
B.1 API Misuse 1.2 1.2 1.8 0.0 0.0 2.4 0.6 2.9 2.1 1.3 2.4 1.0 1.8 1.3 1.8 4.7 6.5 7.8 0.5 1.3 1.0
B.2 Definition Missing 0.0 2.4 0.0 0.6 0.0 5.5 0.0 0.8 0.3 0.0 0.0 0.0 0.0 0.0 1.0 1.2 2.3 3.3 0.5 2.2 1.8
B.3 Incorrect Boundary Condition Check 0.6 2.4 0.6 0.0 0.0 3.0 0.0 2.4 1.3 2.1 1.9 1.5 0.5 1.5 1.8 4.2 5.2 5.7 2.8 4.3 1.7
B.4 Incorrect Argument 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.8 0.0 0.5 1.0 0.0 0.3 37.7 0.7 0.2 2.8 1.5 0.3 1.0
B.5 Minors 3.7 1.8 1.2 1.2 0.6 0.6 0.6 1.1 0.8 0.5 0.3 0.5 0.0 2.3 1.0 1.8 2.7 1.3 1.8 1.7 2.2
B Runtime Bug 5.5 7.9 3.7 1.8 0.6 11.6 1.2 7.1 5.3 4.0 5.0 4.0 2.3 5.3 43.3 12.5 16.8 21.0 7.2 9.8 7.7
C.1 Misunderstanding and Logic Error 29.3 29.9 34.1 31.1 12.8 20.7 17.1 19.0 18.8 19.8 20.9 12.0 15.5 13.8 26.7 48.3 44.5 65.7 31.3 46.5 37.5
C.2 Hallucination 0.0 1.2 0.0 0.0 0.6 0.0 1.2 7.4 7.7 10.3 8.2 1.8 5.8 3.0 5.3 7.5 10.0 0.5 7.0 3.0 4.5
C.3 Input/Output Format Error 0.0 0.0 0.0 0.0 0.0 0.0 1.2 2.6 1.3 1.1 0.5 2.8 0.8 2.8 3.7 2.8 1.2 2.3 0.3 2.7 1.7
C.4 Minors 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.5 0.3 0.0 0.8 0.3 1.7 2.0 2.7 0.7 0.7 2.7 2.0
C Functional Bug 29.3 31.1 34.1 31.1 13.4 20.7 19.5 29.1 27.8 31.7 29.9 16.5 22.8 19.8 37.3 60.7 58.3 69.2 39.3 54.8 45.7
D Ambiguous Problem 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.8 0.8 0.8 0.8 0.8 0.8 0.8
B.2 Definition Missing. Python requires variables and functions must be defined prior to their use in a program However, occasionally omit the definition of commonly used variables or simple functions. As illustrated in the code below, the LLM overlooks the defining variable MOD, which is commonly employed in algorithmic problems.
@@
B.4 Incorrect Argument. LLMs occasionally disregard the specified input format in problem descriptions, resulting in mismatches in the number or type of arguments in the generated code. As shown in the code below, the problem includes two inputs: the first indicates the number of elements and the second represents the elements to be processed. However, the generated code only sets one parameter to receive the elements.
@@
def min_reversals(directions): # lack the first argument

### v2

@@ 3 LLM-Based Code Generation
3.1 Effectiveness of LLMs in Code Generation.
@@
Performance analysis. Within our experimental setup, a generated code is considered correct if it passes compiler checks, runs without errors, and produces the expected outputs for predefined inputs. The correct code must pass all unit test cases in the benchmark, as different unit test cases cover various semantics.
Table 2: Modalities of three commonly used benchmarks and performance of nine popular LLMs in code generation on these benchmarks. Intro. and Comp. denote Introductory and Competition, respectively. Description denotes the token number of the description. LoC denotes lines of code. $C$ denotes Cyclomatic Complexity of code. API denotes the number of API in the code. Cov denotes branch coverage of the test cases. DSC denotes DeepSeekCoder. DSV denotes DeepSeek-V3. DSR denotes DeepSeek-R1.
Dataset Description LoC $C$ API Cov Open-Source Closed-Source
StarCoder-2 DSC Llama-3 Phi-3 DSV DSR GPT-4 GPT-3.5 Claude-3
HumanEval Plus 131.3 11.3 3.1 3.4 99.3% 61.6% 58.5% 51.8% 67.1% 87.8% 93.3% 85.4% 64.6% 79.3%
MBPP Plus 48.7 6.1 2.2 1.8 99.6% 63.2% 66.4% 61.1% 63.2% 75.4% 81.2% 79.2% 73.9% 74.9%
APPS Plus Intro. 307.7 10.3 3.2 5.6 94.8% 42.0% 45.5% 37.5% 13.5% 57.5% 87.5% 72.0% 56.0% 69.5%
Interview 480.3 27.4 6.4 15.1 90.5% 10.5% 20.5% 17.5% 9.0% 53.0% 66.7% 46.5% 25.0% 38.0%
Comp. 554.9 33.7 6.9 18.4 92.9% 2.0% 1.0% 7.5% 1.5% 41.0% 55.0% 36.0% 8.5% 22.0%
Average 447.6 23.8 5.5 13.0 92.7% 18.2% 22.3% 20.8% 8.0% 50.5% 69.8% 51.5% 29.8% 43.2%
Total Average 304.6 17.8 4.4 8.9 96.0% 35.9% 38.4% 35.1% 30.9% 62.6% 71.2% 63.8% 45.6% 56.7%
Based on Table 2, we observe a distinct performance gap across LLM architectures. DeepSeek-R1, incorporating advanced reasoning capabilities, achieves superior performance with 71.2% overall success rate, representing a significant advancement over traditional autoregressive models. Among non-reasoning architectures, both closed-source and open-source models show varied performance, with GPT-4 (63.8%), DeepSeek-V3 (62.6%), and Claude-3 (56.7%). The gap between closed-source and non-reasoning open-source models thus spans roughly 1.2% to 32.9%. Notably, the introduction of reasoning mechanisms in DeepSeek-R1 yields a 7.4% - 40.3% improvement over non-reasoning models, suggesting that architectural innovations targeting explicit reasoning capabilities, rather than mere parameter scaling, constitute a critical factor in enhancing code generation reliability.
Finding 1: The introduction of reasoning capabilities in DeepSeek-R1 yields a 7.4 - 40.3% improvement in code generation performance over traditional models, while a persistent 1.2% - 32.9% performance gap exists between closed-source and open-source implementations. Implication: Software engineering research should prioritize developing reasoning-enhanced architectures and investigating the architectural factors underlying closed-source model superiority, as these represent the most promising pathways to achieving production-grade code generation rather than pursuing incremental improvements to existing approaches.
Dataset analysis. The performance gap between model categories widens dramatically with increasing problem complexity. On the APPS+ Competition subset, the open-source model StarCoder-2 manages only a 2.0% success rate, versus 36.0% for GPT-4 and 55.0% for DeepSeek-R1. This disparity tracks closely with rising complexity, cyclomatic complexity climbs from 3.1 (HumanEval+) to 6.9 (APPS+ Competition), and average API usage from 3.4 to 18.4, yet even the best models suffer severe performance drops. In particular, DeepSeek-R1’s reasoning enhancements yield a 19.0% absolute improvement over GPT-4 on APPS+ Competition, underscoring that non-reasoning LLMs face fundamental limitations on complex code-generation tasks. This systematic decline in pass rates with task difficulty suggests that bugs in LLM-generated code increase predictably as problems become harder, raising critical reliability concerns for production software development.
@@
Figure 1: The differences of code characteristics between code correctly generated by models and canonical solutions on HumanEval+. $C$ denotes cyclomatic complexity.
Finding 2: LLM code generation performance exhibits systematic degradation that correlates exponentially with problem complexity metrics (cyclomatic complexity, API usage, and code length), with even state-of-the-art models dropping significantly when transitioning from simple to competition-level programming tasks. Implication: Production software systems should restrict LLM-based code generation to well-bounded, low-complexity components and implement rigorous validation pipelines, as the predictable increase in bug density with complexity necessitates human oversight for any code exceeding simple algorithmic tasks.
3.2 Factors for LLMs in Code Generation
@@
In this section, we analyze specific characteristics in the generated code, including the number of lines of code, cyclomatic complexity, number of APIs, and number of comment lines.
@@
Code size analysis. We investigated the differences in code characteristics between the correct code and the canonical solution. Figure 1 shows the differences in lines of code, cyclomatic complexity, and API number on HumanEval+. We can observe that the code generated by most models, including reasoning model, typically has fewer lines but slightly higher complexity, while the API number is similar to that of the canonical solutions. Notably, Llama3 tends to generate shorter and more API-heavy code, which is likely due to Llama3 utilizing APIs to replace parts of the complex code implementation. Due to the page limit, we place the experimental results for the remaining benchmarks in the open-source repository [54].
@@
Finding 3: The code generated by the models typically has fewer lines of code but higher code complexity and a similar API number as compared to the canonical solutions. Implication: Optimizing code generation requires not only minimizing lines of code but also managing complexity and employing proper APIs. In practice, balancing among brevity, complexity, and API usage can lead to more maintainable and efficient solutions.
Number of comments analysis. Figure 2 compares the number of comments in correct versus incorrect code generated by various LLMs across multiple benchmarks. Prior work has shown that introducing a chain-of-thought or “thinking” process can improve LLMs’ reasoning ability in both code reasoning and task planning tasks (e.g., [55, 56, 57, 58, 59, 60]). Moreover, inserting comments during code generation can help models articulate their thought process and potentially enhance downstream performance [19, 61].
However, our results indicate a counterintuitive trend: incorrect code often contains more comments than correct code. For instance, in HumanEval+, incorrect solutions, particularly from Claude-3 and DeepSeekCoder, exhibit both a higher volume and greater variability of comments. A similar pattern emerges in APPS+, where multiple models show a higher median number of comments in incorrect code.
Two factors may contribute to this phenomenon. First, complex training examples often come with more comments, which LLMs may replicate when they are uncertain. Second, when unsure of correctness, LLMs might add explanatory comments in an attempt to clarify their logic—yet these extra comments do not necessarily yield higher accuracy. Indeed, even with more detailed commentary, the overall correctness rate remains low.
As for DeepSeek-R1, we observe a distinctly different pattern: the distribution of comments remains remarkably consistent between correct and incorrect code across all benchmarks. Unlike other models that show increased commenting in incorrect solutions, DeepSeek-R1 maintains a stable comment density regardless of solution correctness. This unique behavior can be attributed to its architectural design and training methodology. DeepSeek-R1 employs a reinforcement learning framework that explicitly separates the ”thinking” process from the final code generation. During training, the model learns to internalize its reasoning chain rather than externalizing it through comments. This architectural choice fundamentally differs from traditional approaches where models are encouraged to verbalize their thought process inline with code. Consequently, DeepSeek-R1’s comments serve purely as code documentation rather than as a reasoning mechanism, resulting in consistent comment patterns independent of solution quality.
In the figure, we highlight (in bold) models whose comment distributions differ significantly ($p<0.05$) between correct and incorrect code and remove outliers for clarity. These findings underscore that while comments can reflect a form of “chain of thought,” they do not inherently translate into more accurate code generation.
Finding 4: Most LLMs generate more comments in incorrect code when uncertain, but DeepSeek-R1 maintains consistent comment density regardless of correctness due to its internalized reasoning architecture. Implication: The lack of correlation between comment density and code quality in advanced models like DeepSeek-R1 indicates that future LLM architectures can achieve superior reasoning performance through internalized thought processes rather than explicit chain-of-thought verbalization, challenging current assumptions about the necessity of explanatory comments for effective code generation.
Figure 2: The difference in the comment-to-code-line ratio between the correct and incorrect code generated by LLMs. SC2 denotes StarCoder-2. DC denotes DeepSeekCoder. LL3 denotes Llama-3. CL3 denotes Claude-3. DSV denotes DeepSeek-V3. DSR denotes DeepSeek-R1. The bold label indicates comments in the incorrect code are significantly (i.e., $p<0.05$) higher than in the correct code.
4 Bugs in Code Generated by LLMs
@@
To fully understand the root causes of bugs generated by LLMs, we adopt a two-phase annotation strategy, automated annotation and manual annotation, categorizing 5,741 bugs generated from nine LLMs, identifying three primary bug types and ten secondary bug types (Section 4.1). We further performed a comprehensive bug analysis, including a detailed investigation into the distribution of various bugs, and discussed findings for different bug types (Section 4.2).
4.1 Taxonomy of Code Generation Bugs
4.1.1 Methodology
@@
Inspired by existing work[13, 62], we first annotated bugs in code generation into three primary bug types: Syntax Bug, Runtime Bug, and Functional Bug. These three primary bug types, identified using the Python interpreter’s output, cover all possible bug types. Based on the primary bug types above, we adopted a two-stage process involving script-based automated annotation and manual annotation. The initial script-based annotation aims to efficiently annotate the primary bug types based on the results generated by the Python interpreter, enhancing the manual detailed annotation. The manual annotation involved experts discussing and refining the taxonomy of secondary bug types, ensuring that all bugs were correctly annotated into appropriate secondary types.
@@
In the first stage, the script collected output from the Python interpreter during code execution and used regular expressions for initial bug annotation. Syntax bugs are identified by matching “Syntax Invalid” or “Syntax Error” in the error messages. Runtime bugs are detected by matching the “Traceback” field. Functional bugs are recognized by the “AssertionError” phrase, as assert statements are used to verify outputs in the experiment. The interpreter does not provide detailed error information for “AssertionError”. Therefore, the script utilized GPT-4 to generate three possible root causes along with explanations, aiding in subsequent manual analysis.
These categories align with the standard software testing hierarchy where bugs are typically classified based on when they are detected: at compile-time (syntax), runtime (execution), or through functional testing. The annotation priority ensures systematic coverage without overlap between categories.
Figure 3: Taxonomy of bugs that occurred in code generated by LLMs.
@@
In the second stage, 22 experts with a background in software engineering, reviewed the incorrect code annotated with primary bug types and independently developed a taxonomy for secondary bug types. Then, the experts combined each individual taxonomy and discussed their findings to establish a final taxonomy. Finally, different experts double-checked the annotation to ensure all bugs were correctly annotated. For bugs with disagreements in annotation, all experts convened for a joint discussion until a consensus was reached. Additionally, experts were required to review and revise the annotation of similar bugs to ensure consistency. Notably, because the code generation process is based on “next code token prediction”, the generated code directly influences subsequent code generation. Therefore, we annotated bugs based on the root cause of the first erroneous code in the generated code. Additionally, during the code annotation process, we found that some problems have incomplete or ambiguous functionality descriptions. The models attempt to infer the missing details to complete the code generation, leading to failures in unit tests. We refer to these issues as Ambiguous Problem Description. The two stages collectively took 2,400 person-hours and culminated in the classification of a taxonomy of three primary bug types and ten secondary bug types, as shown in Figure 3. Detailed introductions of each bug type will be discussed in the following sections.
Table 3: Types of bugs introduced during code generation by six open-source LLMs on three widely used benchmarks (i.e., HumanEval+, MBPP+, and APPS+). SC2 denotes StarCoder-2. DSC denotes DeepSeekCoder. Phi3 denotes Phi-3. DSV denotes DeepSeek-V3. DSR denotes DeepSeek-R1. All values are in %.
Bug Types HumanEval Plus MBPP Plus APPS Plus
SC2 DSC Llama3 Phi3 DSV DSR SC2 DSC Llama3 Phi3 DSV DSR SC2 DSC Llama3 Phi3 DSV DSR
A.1 Incomplete Syntax Structure 0.0 0.0 0.0 0.0 0.0 0.0 0.3 0.5 2.4 1.9 0.0 0.0 0.3 3.5 2.7 0.8 1.2 0.0
A.2 Incorrect Indentation 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.2 0.0 0.0 0.0 0.0
A.3 Library Import Error 3.7 2.4 10.4 0.0 0.0 0.0 0.3 0.0 0.8 0.0 0.0 0.0 0.0 0.0 0.5 0.2 0.0 0.5
A Syntax Bug 3.7 2.4 10.4 0.0 0.0 0.0 0.5 0.5 3.2 1.9 0.0 0.0 0.3 3.7 3.2 1.0 1.2 0.5
B.1 API Misuse 1.2 1.2 1.8 0.0 0.0 0.0 2.9 2.1 1.3 2.4 2.7 1.9 1.8 4.7 6.5 7.8 1.2 1.5
B.2 Definition Missing 0.0 2.4 0.0 0.6 0.6 0.0 0.8 0.3 0.0 0.0 1.9 1.3 1.0 1.2 2.3 3.3 4.0 0.8
B.3 Incorrect Boundary Condition Check 0.6 2.4 0.6 0.0 1.2 0.6 2.4 1.3 2.1 1.9 2.1 2.1 1.8 4.2 5.2 5.7 1.0 0.7
B.4 Incorrect Argument 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.8 0.0 0.5 0.0 0.0 37.7 0.7 0.2 2.8 0.3 0.0
B.5 Minors 3.7 1.8 1.2 1.2 0.0 0.0 1.1 0.8 0.5 0.3 0.3 0.3 1.0 1.8 2.7 1.3 0.0 1.7
B Runtime Bug 5.5 7.9 3.7 1.8 1.8 0.6 7.1 5.3 4.0 5.0 6.9 5.6 43.3 12.5 16.8 21.0 6.5 4.7
C.1 Misunderstanding and Logic Error 29.3 29.9 34.1 31.1 9.8 5.5 19.0 18.8 19.8 20.9 16.1 12.4 26.7 48.3 44.5 65.7 33.7 21.5
C.2 Hallucination 0.0 1.2 0.0 0.0 0.0 0.0 7.4 7.7 10.3 8.2 0.0 0.0 5.3 7.5 10.0 0.5 0.5 0.0
C.3 Input/Output Format Error 0.0 0.0 0.0 0.0 0.0 0.6 2.6 1.3 1.1 0.5 0.8 0.0 3.7 2.8 1.2 2.3 3.7 3.0
C.4 Minors 0.0 0.0 0.0 0.0 0.6 0.0 0.0 0.0 0.5 0.3 0.3 0.3 1.7 2.0 2.7 0.7 1.8 0.5
C Functional Bug 29.3 31.1 34.1 31.1 10.4 6.1 29.1 27.8 31.7 29.9 17.2 12.7 37.3 60.7 58.3 69.2 39.7 25.0
D Ambiguous Problem Description 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.8 0.8 0.8 0.8 0.8 0.8
4.1.2 Type A: Syntax Bug.
@@
Syntax bugs violate the grammatical rules of the programming language being used. These bugs are detected by the interpreter when it tries to parse the code before execution. There are three secondary syntax bug types: Incomplete Syntax Structure, Incorrect Indentation, and Library Import Error.
@@
A.1 Incomplete Syntax Structure. An incomplete syntax structure indicates that the generated code includes an open or partially written syntax element that has not been properly completed. This type of bug includes incomplete statements, unmatched parentheses, unclosed quotes, or missing colons.
# unclosed Parentheses
print(" ".join(map(str, range(1, n + 1)))
@@
A.2 Incorrect Indentation Python relies on indentation to define the scope of loops, conditionals, functions, and other code blocks. Incorrect indentation violates the syntax structure, causing syntax errors or inconsistent semantics.
@@
A.3 Library Import Error. Library import allows Python programs to utilize external code, avoiding redundant development. Common import errors include missing import statements and incorrect import levels. In following example, code incorrectly imports all public functions from the heapq library within the function body. However, this operation is only permitted outside the function body.
def xuDRm():
# import * only allowed at module level
from heapq import *
4.1.3 Type B: Runtime Bug.
@@
Runtime Bugs refer to bugs that arise when the code fails to conform to runtime specifications, which are detected during execution. Based on the taxonomy, there are five secondary runtime bugs: API Misuse, Definition Missing, Incorrect Boundary Condition Check, Incorrect Argument, and Minors.
@@
tup.sort() # ’tuple’ object has no attribute sort
Table 4: Types of bugs introduced during code generation by three closed-source LLMs on three widely used benchmarks (i.e., HumanEval+, MBPP+, and APPS+). All values are in %.
@@
GPT-4 GPT-3.5 Claude-3 GPT-4 GPT-3.5 Claude-3 GPT-4 GPT-3.5 Claude-3
A.1 Incomplete Syntax Structure 0.0 0.6 0.0 0.0 0.5 0.0 0.2 1.5 0.2
A.2 Incorrect Indentation 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.2 0.0
A.3 Library Import Error 0.6 2.4 0.0 0.3 0.5 0.0 1.0 2.0 2.5
A Syntax Bug 0.6 3.0 0.0 0.3 1.0 0.0 0.2 4.7 2.7
B.1 API Misuse 0.0 2.4 0.6 2.4 1.0 1.8 0.5 1.3 1.0
B.2 Definition Missing 0.0 5.5 0.0 0.0 0.0 0.0 0.5 2.2 1.8
B.3 Incorrect Boundary Condition Check 0.0 3.0 0.0 1.5 0.5 1.5 2.8 4.3 1.7
B.4 Incorrect Argument 0.0 0.0 0.0 1.0 0.0 0.3 1.5 0.3 1.0
B.5 Minors 0.6 0.6 0.6 0.5 0.0 2.3 1.8 1.7 2.2
B Runtime Bug 0.6 11.6 1.2 4.0 4.0 2.3 7.2 9.8 7.7
C.1 Misunderstanding and Logic Error 12.8 20.7 17.1 12.0 15.5 13.8 31.3 46.5 37.5
C.2 Hallucination 0.6 0.0 1.2 1.8 5.8 3.0 7.0 3.0 4.5
C.3 Input/Output Format Error 0.0 0.0 1.2 2.8 0.8 2.8 0.3 2.7 1.7
C.4 Minors 0.0 0.0 0.0 0.0 0.8 0.3 0.7 2.7 2.0
C Functional Bug 13.4 20.7 19.5 16.5 22.8 19.8 39.3 54.8 45.7
D Ambiguous Problem Description 0.0 0.0 0.0 0.0 0.0 0.0 0.8 0.8 0.8
B.2 Definition Missing. Python requires that variables and functions must be defined prior to their usage in a program. However, LLMs occasionally omit the definition of commonly used variables or simple functions. As illustrated in the code below, the LLM overlooks the defining variable MOD, which is commonly employed in algorithmic problems.
@@
B.4 Incorrect Argument. LLMs occasionally disregard the specified input format in problem descriptions, resulting in mismatches with the number or type of arguments. As shown in the code below, the problem includes two inputs: the first indicates the number of elements, and the second represents the elements to be processed. However, the generated code only sets one function argument to receive the inputs.
----- Input -----
@@
def min_reversal(directions): # lack an argument